Williams Optical Inc. is considering a new lean product cell. The present manufacturing approach produces a product in four separate steps. The production batch sizes are 64 units. The process time for each step is as follows:

Process Step 1 7 minutes
Process Step 2 5 minutes
Process Step 3 4 minutes
Process Step 4 9 minutes

The time required to move each batch between steps is 14 minutes. In addition, the time to move raw materials to Process Step 1 is also 14 minutes, and the time to move completed units from Process Step 4 to finished goods inventory is 14 minutes.

The new lean layout will allow the company to reduce the batch sizes from 64 units to 4 units. The time required to move each batch between steps and the inventory locations will be reduced to 2 minutes. The processing time in each step will stay the same.

Determine the value-added, non-value-added, and total lead times, and the value-added ratio under the present and proposed production approaches. If required, round percentages to one decimal place.
Answer #1

Ans: Present Approach:

1, Value added time:( process time step 1+ process time step 2+ process time step 3+ process time step 4

=(7 mins+ 5 mins+ 4 mins+ 9 mins)

=(25 mins)

2.Non value added= total within batch time + move time

= ( total time required for the 4 steps* (batch size-1)+ move time

= (25 mis*(64-1)+39

=(25*63+39)

=( 1614 min)

3. total lead time= value time+ non value time

=25 mins+ 1614 mins

= (1639 mins)

4. Value added ratio= Value added time/ lead time

= 25 mins/ 1639 mins

= 1.53 mins

Proposed approach:

1. Value added time= ( process time step 1+ process time step 2+ process time step 3+ process time step 4

=(7+5+4+9)

=25 mins

2. Non value added time= total within batch time + move time

=( 7+5+4+9)*(4-1)+28

=(25*3+28)= 103 mins

3 total lead time= ( value added time+ non value added time)

( 103+25)

128 mins

4. Value added ratio= value added time/ total lead time

=25 /128 mins

= ( 19.53 %)